answersLogoWhite

0

drawing a club

total club =13

P(Club)= 13/52

drawing a 7

total 7 = 4

P(7) = 4/52

P(a club and a 7) = P(Club) + P(7) as events are mutualy exculisve

P(a club and a 7) = 13/52+4/52

P(a club and a 7) = 17/52

User Avatar

Wiki User

∙ 15y ago
User Avatar

What else can I help you with?

Related Questions

What is the probability of drawing a heart or club from a pack of cards?

Probability of drawing a heart: 1/4 Probability of drawing a club: 1/4 Probability of drawing a heart or a club: 1/4 + 1/4 = 2/4 = 1/2


What is the probability of drawing a 7 and a club from a deck of cards?

The answer depends on how many cards are drawn.


What is the probability of drawing a club in a pack of playing cards?

In a standard deck of 52 playing cards, there are 13 clubs. The probability of drawing a club is calculated by dividing the number of clubs by the total number of cards. Therefore, the probability of drawing a club is 13/52, which simplifies to 1/4 or 25%.


What is the probability of drawing a club from a deck of 52 playing cards?

There are 13 clubs in a deck of 52 cards. The probability of drawing 1 club from a deck of 52 is 13/52 or 1/4.


What is the probability of drawing a king knowing it is a club?

1/15 actualy its not 1/52 if its a club


Drawing a club from a deck of cards?

There is a 13 in 52, or 1 in 4, or 0.25 probability of drawing a club from a standard deck of 52 cards.


What is probability of drawing a seven or a nine out of cards?

Probability of drawing a 7 or 9 is the probability of draw a 7 plus the probability of draw a 9 which is 4/52 + 4/52 = 8/52 = 2/13 = 0.1538.


Given that you just drew either a spade or a club what is the probability that you also drew a king?

The probability of drawing a king given that you drew a spade or a club is 2 out of 26, or 1 out of 13. This is because there are 2 kings (one from spades and one from clubs) out of a total of 26 spade and club cards.


What is the probability that one is a club and one is a diamond?

To determine the probability that one card drawn is a club and the other is a diamond from a standard deck of 52 cards, you can use the concept of combinations. There are 13 clubs and 13 diamonds in the deck. The probability of drawing one club and one diamond in two draws (without replacement) can be calculated as follows: the probability of drawing a club first and then a diamond is (13/52) * (13/51), and the probability of drawing a diamond first and then a club is (13/52) * (13/51). Adding these two probabilities gives you the total probability of one card being a club and the other a diamond. The final probability is approximately 0.25 or 25%.


What is the probability of drawing a king of club from a deck of 52 cards?

1 out of 52


What is the probability of drawing a gray marble?

4 out of 7


If one card is drawn from a deck of 52 cards what is the probability that it will either be a club or seven of hearts?

In a standard deck of 52 cards, there are 13 clubs and 1 seven of hearts. The probability of drawing a club or the seven of hearts, then, is 14 in 52, or 7 in 26.

Trending Questions
Do girls like big penises? When can the standard error of the estimate be used to construct a prediction interval about a value y'? Which kind of graph would be best for depicting data collected on the weight of a baby every month for six months? What is a Type II Error - a false-negative error? What happens to the place value of a number when you multiply by 10? Which of the following measures of central tendency can be used at any level of measurement? How many outcomes or combinations are possible with three throws of dice? What is linearity error? What is the probability of three people flipping a coin and two having the same side and the other not? How could you plot more than one dependent variable on a line graph? What are the chances a girl is pregnant if the guy ejaculated inside her two weeks after her period? What is ideal average in statistics? What is the probability in picking a vowel in mathematical? Which visual can not be used to compare items over time. A table . A pie chart . A line graph . A bar chart? How many even four digit numbers can be made using 02356 where the first digit is cannot be 0 and no digits are repeated? What is valences in discrete math? What standards of guidelines would you recommend for deciding whether statistical differences between men and women or non minorities and minorities reflect discrimination occurring in an organization? Number of tornadoes in 2010 listed by state? What is sampling with replacement and why is it not used? What is data base sorting?